On a blueprint, the scale indicates that 6 centimeters represent 15 feet. What is the length of a room that is 15.6 centimeters long and 6 centimeters wide on the blueprint?
Answer by checkley77(12569)

6/15.6=15/x
6x=15.6*15
6x=234
x=234/6
x=39 ft. is the answer for the length.